OpenStudy (poohb):

Choose the correct procedure for calculating (f + g)(-2) if f(x) = 13x2 and g(x) = 7x - 5.

OpenStudy (alexwee123):

(f+g)(x)= 13x^2 + 7x-5 (f+g)(2) means to sub in 2 for x for (f+g)(x)

OpenStudy (poohb):

so what would my answer be

OpenStudy (poohb):

Choose the correct procedure for calculating (f - g)(3) if f(x) = 3x - 2 and g(x) = x2 - 1.

OpenStudy (anonymous):

So you would do (3x-2)-(x^2-1)=3x-2-x^2+1=>3x-x^2-1, now substitute 3 for x, we get 3(3)-3^2-1=9-9-1=-1